Ian reports a mild wet morning yesterday, but getting out fine later on. He was on his own, with not a lot to do as the T2/Q6 was back in traffic after its repairs, and the J27 was still stuck at New Bridge because of the congestion at Grosmont caused by the continuing P/W work on the new point for No 6 Road.
However, Charlie Dore wanted to colour code our lifting equipment. So, after opening up Deviation Shed, he came up with his tin of paint and colour coded all our lifting equipment Green, including the lifting jack and chain block. Colour code for the next 6 months is therefore Green. Ian then went down to the MPD and helped out a little, cleaning two of Repton's valve covers. Repton at the moment is having its valves removed and carbon cleaned from the heads and rings. At the end of the day, the middle valve was still to extract after many hours moving it to a point where it just wouldn’t move any farther forward. Hopefully it will have been removed today!
Other things going on at the MPD was the painting of the 9F No 92134 and tender in black (with the later BR emblem) being done by Heritage Painting. Locomotives working yesterday were B1 No 1264 , Black Five No 5428 Eric Treacy, and T2 No 2238.
Meanwhile, Bill Dobson and Gordon Wells spent the day at the NRM in York looking for safety valve drawings for the T2/Q6 to enable the new castings to be machined, but didn’t have any success. However, further information has come to light which will hopefully lead to finding the required drawings, and I have been charged with trying to track them down over the next week or so. Bill is also anxious to find drawings for the T2 piston, both head and rod. Will keep me out of mischief!